# Nearsightedness in Materials with Indirect Band Gap

### v1

@@ 6. Proofs of the Main Results
$\begin{split}\left|\frac{\partial\big{[}\rho(u)\big{]}_{\ell\ell,ab}}{\partial[u(k)]_{i_{1}}}\right|&=\left|\oint_{\mathscr{C}}\left[\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}\,\frac{\partial{\mathcal{H}}(u)}{\partial[u(k)]_{i}}\,\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}\right]_{\ell\ell,ab}\frac{{\,{\rm d}}z}{2\pi i}\right|\\ &\leq C\left|\left[\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}\frac{\partial[{\mathcal{H}}(u)]}{\partial[u(k)]_{i_{1}}}(z-{\mathcal{H}}(u))^{-1}\right]_{\ell\ell,ab}\right|\\ &\leq\sum_{m,n\in\Lambda}\left(\frac{C_{q}(\mathfrak{d}^{\mathrm{ref}})}{(1+r_{\ell m})^{q}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})r_{\ell m}}+C(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})(|\ell|+|m|)}\right)e^{-\gamma_{0}\mathfrak{m}(r_{mk}+r_{kn})}\\ &\hskip 36.98866pt\cdot\left(\frac{C_{q}(\mathfrak{d}^{\mathrm{ref}})}{(1+r_{n\ell})^{q}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})r_{n\ell}}+C(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})(|n|+|\ell|)}\right)\\ &\leq\sum_{m\in\Lambda}\left[C_{q}(\mathfrak{d}^{\mathrm{ref}})\right]^{2}\frac{e^{-2[\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})r_{\ell m}+\gamma_{0}\mathfrak{m}r_{mk}]}}{(1+r_{\ell m})^{2q}}+[C(u)]^{2}\sum_{m\in\Lambda}e^{-[\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})(|\ell|+|m|)+2\gamma_{0}\mathfrak{m}r_{mk}]}\\ &\hskip 36.98866pt+C(u)C_{q}(\mathfrak{d}^{\mathrm{ref}})\sum_{m\in\Lambda}\frac{e^{-[\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})r_{\ell m}+\gamma_{0}\mathfrak{m}r_{mk}]}}{(1+r_{\ell m})^{q}}\sum_{n\in\Lambda}e^{-[\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})(|\ell|+|n|)+\gamma_{0}\mathfrak{m}r_{kn}]}.\end{split}$ (6.44)
Note that the following estimates: for fixed $\ell,k\in\Lambda,\gamma\geq\beta>0$, we have
$\sum_{m\in\Lambda}\frac{e^{-(\beta r_{\ell m}+\gamma r_{mk})}}{(1+r_{\ell m})^{q}}$ $\leq\sum_{m\in\Lambda}\frac{e^{-(\frac{\beta}{2}r_{\ell k}+\frac{\gamma}{2}r_{mk})}}{(1+|r_{\ell k}-r_{mk}|)^{q}}$
$\leq Ce^{-\frac{\beta}{2}r_{\ell k}}\left(\int_{0}^{r_{\ell k}}\frac{e^{-\frac{\gamma}{2}r}r^{d-1}}{(1+|r_{\ell k}-r|)^{q}}{\,{\rm d}}r+\int_{r_{\ell k}}^{\infty}\frac{e^{-\frac{\gamma}{2}r}r^{d-1}}{(1+|r_{\ell k}-r|)^{q}}{\,{\rm d}}r\right)$
$\leq C{e^{-{\beta}r_{\ell k}}}\left(-\int_{0}^{r_{\ell k}}\frac{e^{\frac{\gamma}{2}r}(r_{\ell k}-r)^{d-1}}{(1+r)^{q}}{\,{\rm d}}r+\int_{0}^{\infty}\frac{e^{-\frac{\gamma}{2}r}(r_{\ell k}+r)^{d-1}}{(1+r)^{q}}{\,{\rm d}}r\right)$
$\leq C\frac{e^{-\beta r_{\ell k}}}{(1+r_{\ell k})^{q}},$ (6.45)
and
$\sum_{m\in\Lambda}e^{-(\beta|m|+\gamma r_{mk})}\leq e^{-\tfrac{\beta}{2}|k|}\Big{(}\sum_{m\in\Lambda}e^{-\tfrac{\gamma}{2}r_{mk}}\Big{)}\leq Ce^{-\frac{\beta}{2}|k|},$ (6.46)
for some constant $C>0$ depending on $d,\mathfrak{m},\gamma$, then we can bound (6.44) by
@@
$\begin{split}\left|\frac{\partial\big{[}\rho(u)\big{]}_{\ell\ell,ab}}{\partial[u(k)]_{i_{1}}}\right|&\leq\frac{C^{2}_{q}(\mathsf{gap}_{-})}{(1+r_{\ell k})^{2q}}e^{-2\eta_{-}r_{\ell k}}+[C(u)]^{2}e^{-2\eta_{-}(|\ell|+|k|)}+\frac{C(u)C_{q}(\mathsf{gap}_{-})}{(1+r_{\ell k})^{q}}e^{-\eta_{-}(|\ell|+|k|+r_{\ell k})}\\ &\leq\Big{[}\frac{C_{q}(\mathsf{gap}_{-})}{(1+r_{\ell k})^{q}}+C(u)\Big{]}^{2}e^{-2\eta_{-}r_{\ell k}},\end{split}$
which completes the proof. ∎
@@ Appendix A Bloch Transform for the Tight Binding Hamiltonian
$\langle\hat{\psi},\hat{\phi}\rangle_{L^{2}\left(\Gamma^{*},\ell^{2}(\Lambda_{0}\times\Xi)\right)}\coloneqq\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{\Gamma^{*}}\langle\hat{\psi}_{\bm{\xi}}(\cdot),\hat{\phi}_{\bm{\xi}}(\cdot)\rangle_{\ell^{2}(\Lambda_{0}\times\Xi)}{\,{\rm d}}\bm{\xi}=\sum_{\begin{subarray}{c}k_{0}\in\Lambda_{0},a\in\Xi\end{subarray}}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{\Gamma^{*}}\hat{\psi}^{*}_{\bm{\xi}}(k_{0};a)\hat{\phi}_{\bm{\xi}}(k_{0};a){\,{\rm d}}\bm{\xi}.$
@@
$\begin{split}(U^{*}\hat{\psi}_{\bm{\xi}})(k_{0}+\mathsf{A}\alpha):=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{\Gamma^{*}}\hat{\psi}_{\bm{\xi}}(k_{0};a)e^{i\bm{\xi}\cdot(k_{0}+\mathsf{A}\alpha)}{\,{\rm d}}\bm{\xi},\end{split}$
@@
$U{\mathcal{H}}^{\mathrm{ref}}U^{*}=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{\Gamma^{*}}^{\oplus}{\mathcal{H}}_{\boldsymbol{\xi}}{\,{\rm d}}\bm{\xi},$
@@
$[{\mathcal{H}}_{\boldsymbol{\xi}}]_{\ell_{0}k_{0}}^{ab}\coloneqq\sum_{\alpha\in\mathbb{Z}^{d}}[{\mathcal{H}}^{\mathrm{ref}}]^{ab}_{\ell_{0}+\mathsf{A}\alpha,k_{0}}e^{-i\bm{\xi}\cdot(\ell_{0}-k_{0}+\mathsf{A}\alpha)}\quad\quad\forall\bm{\xi}\in\Gamma^{*},\,\ell_{0},k_{0}\in\Lambda_{0},$
@@
${\mathcal{H}}_{\boldsymbol{\xi}}u_{n}(\bm{\xi})=\varepsilon_{n}(\bm{\xi})u_{n}(\bm{\xi}),\quad n=1,\dots,MN_{\mathrm{b}},$ (A.1)

### v2

@@ 6.3. Proof of the Theorem 3
(6.32) $\begin{split}\left\|{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}\right\|_{\mathrm{F}}^{2}=&\sum_{\begin{subarray}{c}\ell_{0},k_{0}\in\Lambda_{0}\\ 1\leq a,b\leq N_{\mathrm{b}}\end{subarray}}\left|\big[{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}\big]_{\ell_{0}k_{0}}^{ab}\right|^{2}\leq(N_{\mathrm{b}}h_{0})^{2}\sum_{\begin{subarray}{c}\ell_{0},k_{0}\in\Lambda_{0}\end{subarray}}\left(\sum_{\alpha\in\mathbb{Z}^{d}}e^{-\gamma_{0}|\ell_{0}-k_{0}+\mathsf{A}\alpha|}e^{|\ell_{0}-k_{0}+\mathsf{A}\alpha||\bm{\zeta}|}\right)^{2}\\ \leq&(MN_{\mathrm{b}}h_{0})^{2}\sup_{\begin{subarray}{c}\ell_{0},k_{0}\in\Lambda_{0}\end{subarray}}\left(\sum_{\alpha\in\mathbb{Z}^{d}}e^{-(\gamma_{0}-|\bm{\zeta}|)|\ell_{0}-k_{0}+\mathsf{A}\alpha|}\right)^{2}\\ \leq&\left(C_{\mathfrak{m},d}MN_{\mathrm{b}}h_{0}(\gamma_{0}-|\bm{\zeta}|)^{-d}\right)^{2}.\end{split}$
By choosing an appropriate radius specifically for $|\bm{\zeta}|\leq\tfrac{1}{2}\min\{\gamma_{0},\delta_{*}\}$, we have $\left\|{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}\right\|_{\mathrm{F}}<\infty$. Furthermore, from (6.31), for $n\in\{N_{0},N_{0}+1\}$, it follows that
(6.33) $|\operatorname{Re}\varepsilon_{n}(\bm{\xi}+i\bm{\zeta})-\varepsilon_{n}(\bm{\xi})|\leq\frac{1}{2}\sup_{{\bm{\omega}}\in\mathbb{C}_{\delta_{*}/2}^{d}\cap\mathbb{C}_{\gamma_{0}/2}^{d}}\|\nabla^{2}\varepsilon_{n}({\bm{\omega}})\|_{\rm F}|\bm{\zeta}|^{2}\leq m_{\infty}|\bm{\zeta}|^{2}.$
For any $\bm{\xi},\bm{\zeta}\in\mathbb{R}^{d}$ and $\ell,k\in\Lambda$, we denote $\begin{pmatrix}\bm{\xi}\\ \bm{\zeta}\end{pmatrix}:\begin{pmatrix}\ell\\ k\end{pmatrix}\coloneqq\bm{\xi}\cdot\ell+\bm{\zeta}\cdot k$ for brevity. Following the proof of Theorem 3, we choose a suitable contour $\mathscr{C}\supset i[-R,R]$ for some $R>0$ to be determined later, and write
$\frac{\partial\rho_{\ell\ell}}{\partial[\bm{r}_{k}]_{j}}$ $\coloneqq\frac{\partial\big[\rho(u)\big]_{\ell\ell}}{\partial[u(k)]_{j}}\Bigg|_{u=\mathbf{0}}$
$=\oint_{\mathscr{C}+\varepsilon_{\rm F}}\left[\big(z-{\mathcal{H}}^{\rm ref}\big)^{-1}\,\frac{\partial{\mathcal{H}}(u)}{\partial[u(k)]_{j}}\bigg|_{u=\mathbf{0}}\,\big(z-{\mathcal{H}}^{\rm ref}\big)^{-1}\right]_{\ell\ell}\frac{{\,{\rm d}}z}{2\pi i}$
$=\sum_{mn}\frac{\partial{\mathcal{H}}_{mn}}{\partial[\bm{r}_{k}]_{j}}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{(\Gamma^{*})^{2}}\left[\oint_{\mathscr{C}+\nu(\bm{\xi}_{1},\bm{\xi}_{2})}\big(z-{\mathcal{H}}_{\bm{\xi}_{1}}\big)^{-1}_{\ell_{0}m_{0}}\big(z-{\mathcal{H}}_{\bm{\xi}_{2}}\big)^{-1}_{n_{0}\ell_{0}}\frac{{\,{\rm d}}z}{2\pi i}\right]e^{i\binom{\bm{\xi}_{1}}{\bm{\xi}_{2}}:\binom{\ell-m}{n-\ell}}{\,{\rm d}}\bm{\xi}_{1}{\,{\rm d}}\bm{\xi}_{2}$
$\eqqcolon\sum_{mn}\frac{\partial{\mathcal{H}}_{mn}}{\partial[\bm{r}_{k}]_{j}}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{(\Gamma^{*})^{2}}\widehat{G}(\bm{\xi}_{1},\bm{\xi}_{2})_{\ell_{0}m_{0},n_{0}\ell_{0}}e^{i\binom{\bm{\xi}_{1}}{\bm{\xi}_{2}}:\binom{\ell-m}{n-\ell}}{\,{\rm d}}\bm{\xi}_{1}{\,{\rm d}}\bm{\xi}_{2}$
(6.34) $\eqqcolon\sum_{mn}\frac{\partial{\mathcal{H}}_{mn}}{\partial[\bm{r}_{k}]_{j}}\,G_{\ell m,n\ell},$
For any $1\leq a,b\leq N_{\mathrm{b}}$, we write $\frac{\partial\rho_{\ell\ell,ab}}{\partial[\bm{r}_{k}]_{j}}=\sum_{mn,a^{\prime}b^{\prime}}\frac{\partial{\mathcal{H}}_{mn,a^{\prime}b^{\prime}}}{\partial[\bm{r}_{k}]_{j}}\,G_{\ell m,n\ell}^{aa^{\prime},bb^{\prime}}$. The decay properties of (6.3) relies on the analyticity of the product function defined by
(6.35) $\widehat{G}(\bm{\xi}_{1},\bm{\xi}_{2})\coloneqq\oint_{\mathscr{C}+\nu(\bm{\xi}_{1},\bm{\xi}_{2})}\big(z-{\mathcal{H}}_{\bm{\xi}_{1}}\big)^{-1}\otimes\big(z-{\mathcal{H}}_{\bm{\xi}_{2}}\big)^{-1}\frac{{\,{\rm d}}z}{2\pi i},\quad(\bm{\xi}_{1},\bm{\xi}_{2})\in(\Gamma^{*})^{2},$
where the contour shift satisfies $\max\{\varepsilon_{N_{0}}(\bm{\xi}_{1}),\varepsilon_{N_{0}}(\bm{\xi}_{2})\}<\nu(\bm{\xi}_{1},\bm{\xi}_{2})<\min\{\varepsilon_{N_{0}+1}(\bm{\xi}_{1}),\varepsilon_{N_{0}+1}(\bm{\xi}_{2})\}$ for all $(\bm{\xi}_{1},\bm{\xi}_{2})\in(\Gamma^{*})^{2}$. Consequently, the analyticity of $\widehat{G}$ depends on the regularity of the mapping $\nu:(\Gamma^{*})^{2}\to\mathbb{C}$. Analogously to the proof of Section 3, our goal is to construct a continuous shift $\nu$ and identify its domain of analyticity.
First, as in the proof of Section 3, we define
(6.36) $D_{\rm R}\coloneqq\left\{\bm{\xi}+i\bm{\zeta}\colon\bm{\xi}\in\Gamma^{*},|\bm{\zeta}|\leq\min\Big\{\tfrac{1}{2}\gamma_{0},\tfrac{1}{2}\delta_{*},\sqrt{\tfrac{\alpha}{m_{\infty}}\mathsf{gap}_{-}}\Big\}\right\}$
with $\alpha:=\frac{1}{4}$. For ${\bm{\omega}_{l}}\in D_{\rm R}$ for $l=1,2$, we define
$\nu_{0}({\bm{\omega}_{1}},{\bm{\omega}_{2}})\coloneqq\frac{\min_{\bm{\xi}_{1},\bm{\xi}_{2}}\varepsilon_{N_{0}+1}+\max_{\bm{\xi}_{1},\bm{\xi}_{2}}\varepsilon_{N_{0}}}{2}\in\mathbb{R}$
and construct a polynomial approximation $\nu$ satisfying
(6.37) $|\operatorname{Re}\nu({\bm{\omega}_{1}},{\bm{\omega}_{2}})-\nu_{0}({\bm{\omega}_{1}},{\bm{\omega}_{2}})|\leq\frac{\alpha}{2}\mathsf{gap}_{-},\quad({\bm{\omega}_{1}},{\bm{\omega}_{2}})\in(D_{\rm R})^{2}.$
In the exact same way as in (6.10), we have
${\rm dist}\left(\operatorname{Re}\nu({\bm{\omega}}_{1},{\bm{\omega}}_{2}),\operatorname{Re}\sigma({\mathcal{H}}_{{\bm{\omega}}_{1}})\cup\operatorname{Re}\sigma({\mathcal{H}}_{{\bm{\omega}}_{2}})\right)\geq\frac{1}{4}\mathsf{gap}_{-},$
for all ${\bm{\omega}}_{1},\bm{\omega}_{2}\in D_{\rm R}$. In particular, $\left(\mathscr{C}+\nu({\bm{\omega}}_{1},{\bm{\omega}}_{2})\right)\cap[\sigma({\mathcal{H}}_{{\bm{\omega}}_{1}})\cup\sigma({\mathcal{H}}_{{\bm{\omega}}_{2}})]=\emptyset$ for all $\bm{\omega}_{1},\bm{\omega}_{2}\in D_{\rm R}$. As a result, the function $\widehat{G}$ (6.35) is analytic on $(D_{\rm R})^{2}$.
Let us define the decay rate $\widetilde{\eta}_{-}\coloneqq\min\big\{\tfrac{1}{2}\gamma_{0},\tfrac{1}{2}\delta_{*},\sqrt{\tfrac{\alpha}{m_{\infty}}\mathsf{gap}_{-}}\big\}$. Then, for any ${\bm{\omega}}_{l}=\bm{\xi}_{l}+i\bm{\zeta}_{l}$ with $|\bm{\zeta}_{l}|\leq\widetilde{\eta}_{-}$, $l=1,2$, we have
$\tilde{C}^{2}\coloneqq\int_{(\Gamma^{*})^{2}}\big|\widehat{G}(\bm{\omega}_{1},\bm{\omega}_{2})_{\ell_{0}m_{0},n_{0}\ell_{0}}^{aa^{\prime},bb^{\prime}}\big|^{2}{\,{\rm d}}\bm{\xi}_{1}{\,{\rm d}}\bm{\xi}_{2}=\sum_{mn}\left|G_{\ell m,n\ell}^{aa^{\prime},bb^{\prime}}\right|^{2}e^{2\,\bm{\zeta}_{1}\cdot(\ell-m)+2\,\bm{\zeta}_{2}\cdot(n-\ell)}<\infty,$
Note that $\big|\widehat{G}(\bm{\omega}_{1},\bm{\omega}_{2})_{\ell_{0}m_{0},n_{0}\ell_{0}}^{aa^{\prime},bb^{\prime}}\big|\leq(\mathsf{gap}_{-})^{-2}$ for any $(\bm{\omega}_{1},\bm{\omega}_{2})\in(D_{\rm R})^{2}$, so $|G_{\ell m,n\ell}^{aa^{\prime},bb^{\prime}}|\leq\frac{\tilde{C}}{\mathsf{gap}_{-}^{2}}e^{-\widetilde{\eta}_{-}(r_{\ell m}+r_{n\ell})}$. Combining this with (6.3), we obtain
@@
$\left|\frac{\partial\rho_{\ell\ell,ab}}{\partial[\bm{r}_{k}]_{j}}\right|$ $\leq\frac{1}{2\pi}\ell(\mathscr{C})h_{0}(MN_{\mathrm{b}})^{2}\frac{\tilde{C}}{\mathsf{gap}_{-}^{2}}\sum_{mn}\left|\frac{\partial{\mathcal{H}}_{mn}^{a^{\prime}b^{\prime}}}{\partial[\bm{r}_{k}]_{j}}\right|\left|G_{\ell m,n\ell}^{aa^{\prime},bb^{\prime}}\right|$
$\leq\frac{1}{2\pi}\ell(\mathscr{C})h_{0}(MN_{\mathrm{b}})^{2}\frac{\tilde{C}}{\mathsf{gap}_{-}^{2}}\sum_{mn}e^{-\gamma_{0}(r_{mk}+r_{kn})}e^{-\widetilde{\eta}_{-}(r_{\ell m}+r_{\ell n})}$
$\leq\frac{\tilde{C}}{\mathsf{gap}_{-}^{2}}e^{-\widetilde{\eta}_{-}r_{\ell k}},$
where the constant $\tilde{C}>0$ depends on $h_{0},\gamma_{0},d,\mathfrak{m},M$, and $N_{\mathrm{b}}$.
For the nearsightedness estimate (3.12), define $\widetilde{\eta}_{+}\coloneqq\min\big\{\tfrac{1}{2}\gamma_{0},\tfrac{1}{2}\delta_{*},\sqrt{\tfrac{\alpha}{m_{\infty}}\mathsf{gap}_{+}}\big\}$. Together with the perturbation estimate (6.33), the same argument as in the proof of Theorem 3 (with $\eta_{+}$ replaced by $\widetilde{\eta}_{+}$), shows that (6.10) holds for all $|\bm{\xi}|<\widetilde{\eta}_{+}$ and hence (3.12) follows.
@@ Appendix A Bloch Transform for the Tight Binding Hamiltonian
$\langle\hat{\psi},\hat{\phi}\rangle_{L^{2}\left(\Gamma^{*},\ell^{2}(\Lambda_{0}\times\Xi)\right)}\coloneqq\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{\Gamma^{*}}\langle\hat{\psi}_{\bm{\xi}}(\cdot),\hat{\phi}_{\bm{\xi}}(\cdot)\rangle_{\ell^{2}(\Lambda_{0}\times\Xi)}{\,{\rm d}}\bm{\xi}=\sum_{\begin{subarray}{c}k_{0}\in\Lambda_{0},a\in\Xi\end{subarray}}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{\Gamma^{*}}\hat{\psi}^{*}_{\bm{\xi}}(k_{0};a)\hat{\phi}_{\bm{\xi}}(k_{0};a){\,{\rm d}}\bm{\xi}.$
@@
$\begin{split}(U^{*}\hat{\psi}_{\bm{\xi}})(k_{0}+\mathsf{A}\alpha):=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{\Gamma^{*}}\hat{\psi}_{\bm{\xi}}(k_{0};a)e^{i\bm{\xi}\cdot(k_{0}+\mathsf{A}\alpha)}{\,{\rm d}}\bm{\xi},\end{split}$
@@
$U{\mathcal{H}}^{\mathrm{ref}}U^{*}=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{\Gamma^{*}}^{\oplus}{\mathcal{H}}_{\boldsymbol{\xi}}{\,{\rm d}}\bm{\xi},$
@@
$[{\mathcal{H}}_{\boldsymbol{\xi}}]_{\ell_{0}k_{0}}^{ab}\coloneqq\sum_{\alpha\in\mathbb{Z}^{d}}[{\mathcal{H}}^{\mathrm{ref}}]^{ab}_{\ell_{0}+\mathsf{A}\alpha,k_{0}}e^{-i\bm{\xi}\cdot(\ell_{0}-k_{0}+\mathsf{A}\alpha)},\quad\forall\bm{\xi}\in\Gamma^{*},\,\ell_{0},k_{0}\in\Lambda_{0},$
@@
(A.1) ${\mathcal{H}}_{\boldsymbol{\xi}}u_{n}(\bm{\xi})=\varepsilon_{n}(\bm{\xi})u_{n}(\bm{\xi}),\quad n=1,\dots,MN_{\mathrm{b}},$